# Gaussian Scenes: Pose-Free Sparse-View Scene Reconstruction using Depth-Enhanced Diffusion Priors

### v2

@@ Abstract
In this work, we introduce a generative approach for pose-free (without camera parameters) reconstruction of $360^{\circ}$scenes from a sparse set of 2D images. Pose-free scene reconstruction from incomplete, pose-free observations is usually regularized with depth estimation or 3D foundational priors. While recent advances have enabled sparse-view reconstruction of large complex scenes (with high degree of foreground and background detail) with known camera poses using view-conditioned generative priors, these methods cannot be directly adapted for the pose-free setting when ground-truth poses are not available during evaluation. To address this, we propose an image-to-image generative model designed to inpaint missing details and remove artifacts in novel view renders and depth maps of a 3D scene. We introduce context and geometry conditioning using Feature-wise Linear Modulation (FiLM) modulation layers as a lightweight alternative to cross-attention and also propose a novel confidence measure for 3D Gaussian splat representations to allow for better detection of these artifacts. By progressively integrating these novel views in a Gaussian-SLAM-inspired process, we achieve a multi-view-consistent 3D representation. Evaluations on the MipNeRF360 and DL3DV-10K benchmark dataset demonstrate that our method surpasses existing pose-free techniques and performs competitively with state-of-the-art posed (precomputed camera parameters are given) reconstruction methods in complex $360^{\circ}$scenes.
@@ 1 Introduction
Although most previous sparse view 3D scene reconstruction works are posed and require ground-truth camera poses (DSNeRF (Deng et al., 2022), DNGaussian (Li et al., 2024), SparseGS (Xiong et al., 2023), SparseNeRF (Wang et al., 2023a), DietNeRF (Jain et al., 2021), FreeNeRF (Yang et al., 2023), RegNeRF (Niemeyer et al., 2022), DiffusioNeRF (Wynn & Turmukhambetov, 2023)), recent works like COGS (Jiang et al., 2024) and InstantSplat (Fan et al., 2024) have found initial success in the pose-free sparse view reconstruction problem. Although they have been successful in reducing 3D reconstruction-related artifacts such as blur, floaters, color, and streak-like artifacts, they lack generative capabilities for a complete $360^{\circ}$reconstruction, which requires robust priors from powerful generative models (Saharia et al., 2022; Rombach et al., 2022). In addition, most of these methods lack mechanisms to robustly identify and locate these artifacts in the reconstructed scenes. Recent methods like ZeroNVS (Sargent et al., 2024), Reconfusion (Wu et al., 2024), and CAT3D (Gao* et al., 2024) incorporate 3D view conditioning for realistic extrapolation but depend on accurate poses and cannot be trivially extended to the pose-free problem. Gaussian Object (Yang et al., 2024), iFusion (Wu et al., 2023), and UpFusion (Nagoor Kani et al., 2024) do provide pose-free generative solutions but remain limited to only object reconstruction rather than scene reconstruction. Other generative solutions like ZeroNVS, ReconFusion, and CAT3D use stronger priors over regularization-based techniques like FreeNeRF, RegNeRF, and DietNeRF but rely on large-scale 3D datasets, video datasets, and compute resources not available to the average researcher. Furthermore, works like ReconFusion and CAT3D remain closed-source with no access to their models, data or other reproducibility data, which limits their usage for the community. We summarize such existing methods and their use cases and attributes in Fig 2.
@@ 2 Method
This section begins with an overview of our method in Sec 2.1, detailing our approach for reconstructing a 3D scene from a sparse set of uncalibrated 2D images. In Sec 2.1, we describe how we initialize a Gaussian point cloud using MASt3R and 3DGS to provide a coarse 3D representation. Sec 2.2 introduces our RGBD image-to-image generative model, which refines rendered novel views by correcting artifacts and filling in missing details. In Sec 2.3, we propose a confidence measure based on cumulative transmittance and Gaussian density to guide the generative diffusion model in identifying unreliable regions in novel-view renders. Sec 2.4 outlines our synthetic dataset creation process, which enables training the generative model with high-quality RGBD supervision. Sec 2.5 details our depth-augmented autoencoder finetuning process to improve latent-space encoding for RGBD data. Sec 2.5.1 explains the fine-tuning of our UNet with synthesized training data to generate photorealistic and geometrically consistent novel views. In Sec 2.5.2, we describe the inference process of our generative model, where novel views are synthesized using RGBD renders and confidence maps. Sec 2.6 presents our iterative 3D Gaussian optimization strategy that progressively integrates novel view constraints into the scene representation. Finally, Sec 2.7 describes a test-time pose alignment step that refines camera poses to align the rendered image with a given test view before evaluation.
@@ 2.1 Algorithm Overview
Given a set of $M$ images =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{M}\}$ of an underlying 3D scene with unknown intrinsics and extrinsics, our goal is to reconstruct the 3D scene, estimate the camera poses of a monocular camera at the $M$ training views, and synthesize novel views at evaluation time given by $N$ unseen test images $\{I_{M+1},I_{M+2},\dots,I_{M+N}\}$.
@@
1:Sparse image set =$\mathcal{I}=\{I_{1},\dots,I_{M}\}$, densification interval $k$, iterations $N$
@@
$\hat{\mathcal{I}}\leftarrow\mathcal{I}$triangleright Initialize training images
@@
8:         $(\hat{I},\hat{D})\leftarrow R_{\pi_{novel}}(\mathcal{G})$triangleright Render image and depth map
@@
10:         Extract semantic context & geometry embeddings $\{c_{clip},c_{geo}\}$triangleright Sec 2.2.2
11:         $(I_{ref},D_{ref})\leftarrow\text{Refine}(\hat{I},\hat{D},\mathcal{C},c_{clip},c_{geo})$triangleright Refine $\hat{I}$, $\hat{D}$ with generative priors (Sec 2.5.2)
12:         ^^$\hat{\mathcal{I}}\leftarrow\hat{\mathcal{I}}\cup\{I_{ref}\}$triangleright Synthesized novel view added as training sample
@@
16:     Optimize 3D Gaussians $\mathcal{G}$triangleright Eq. 6 or 3DGS loss
@@ 2.2.1 Generative Model Architecture
Our model comprises of (1) A variational autoencoder (encoder $\mathcal{E}$, decoder $\mathcal{D}$) operating in a compressed latent space. (2) A UNet denoiser $\epsilon_{\theta}$ predicting noise in diffused latent $z_{t}$ (3) A multi-modal conditioning incorporating RGBD renders, confidence maps, semantic context, and geometric information.
@@ 2.2.2 Multi-modal Conditioning
We initialize $\epsilon_{\theta}$ with pretrained Stable-Diffusion-2 model weights and expand the first convolutional layer to accept additional inputs by concatenating the noisy latent $z_{t}$, the encoded RGBD image $\mathcal{E}(\hat{I})$, the confidence-weighted encoded image $\mathcal{E}(\hat{I}\cdot\mathcal{C})$ and the downsampled confidence map $\hat{\mathcal{C}}$. To ensure view coherence and geometric consistency, we incorporate following conditioning signals:
@@ 2.2.3 Parameter-Efficient FiLM Conditioning
Our FiLM-based approach requires only 8.14M parameters (7.38M for CLIP features, 758K for pose embeddings) compared to 29.8M for an equivalent cross-attention implementation—a 3$\times$ reduction while maintaining comparable quality, enabling more efficient training and faster inference during iterative reconstruction.
@@ 2.3 Pixel-Aligned Confidence Map
Figure 7: Confidence Measure comparison with 3DGS-Enhancer (Liu et al., 2024). Our confidence map accurately identifies artifacts and 00-Gaussian regions in the sparse-view (darker pixels) while  Liu et al. (2024) incorrectly attributes high confidence to regions with overlap of small-scale Gaussians. NVS render from a densely fitted 3DGS representation is provided for reference.
@@ 2.4 RGBD Dataset Creation
For training the additional weights in $\bm{\epsilon_{\theta}}$ for RGBD image-to-image diffusion, we rely on a set $\mathcal{X}=\{(I^{i},\hat{I}^{i},\mathcal{C}^{i},c_{clip}^{i},c_{geo}^{i})_{i=1}^{N}\}$, each containing a clean RGBD image $I^{i}$, an RGBD image with artifacts $\hat{I}^{i}$ and the corresponding confidence map $\mathcal{C}^{i}$, CLIP features of source images $c_{clip}^{i}\in\mathcal{R}^{M\times 768}$, and plücker embeddings of source and target cameras $c_{geo}^{i}\in\mathcal{R}^{(M+1)\times 78}$, to “teach” the diffusion model how to inpaint missing details and detect Gaussian artifacts guided by the confidence map, context and geometry features and generate a clean version of the conditioning image. For this, we build a dataset generation pipeline comprising a high-quality 3DGS model fitted to dense views, a low-quality 3DGS model fitted to few views, and camera interpolation and perturbation modules to use supervision of the high-quality model at viewpoints beyond ground truth camera poses. For a given scene, we fit sparse models for $M\in\{3,6,9,18\}$ number of views. We render $I^{i}$ using the high-quality model and $\hat{I}^{i}$, $\mathcal{C}^{i}$ using the low-quality model. We save the CLIP features $c_{clip}$ and plücker embeddings of the $M$ sparse views and 1 target view per sample for conditioning $\bm{\epsilon_{\theta}}$ Our fine-tuning setup is illustrated in Fig  12 (Sec C).
@@ 2.6 Scene Reconstruction with Generative Priors
Our generative diffusion model provides a generative prior to infer plausible detail in unobserved regions. Despite view conditioning using pose embeddings, the generated images at novel poses lack complete 3D consistency. For this, we devise an iterative strategy where we first sample novel views along an elliptical trajectory fitted to the training views. We initialize the Gaussian optimization with the set of Gaussians $\mathcal{G}$ fitted to the training views (Sec 2.1). Each novel view is added to the training stack at the beginning of every densification step to encourage the optimization to adjust to the distilled scene priors. At every iteration, we sample either an observed or unobserved viewpoint from the current training stack. We bring back Adaptive Density Control to encourage densification of Gaussians in 0-Gaussian regions. We employ the 3DGS objective for the training views. For novel views, we employ the SparseFusion (Zhou & Tulsiani, 2023) objective in the RGB space and a PCC loss for the rendered and denoised depths.
@@ 2.7 Test-time Pose Alignment
[…]
@@
[…]
@@ 3.3 Comparative Results
[…]
@@
[…]
@@ 4 Conclusion
In this work, we present GScenes where we integrate an image-to-image RGBD diffusion model with a pose-free reconstruction pipeline in MASt3R to reconstruct a $360^{\circ}$3D scene from a few uncalibrated 2D images. We introduce context and geometry conditioning through FiLM layers achieving similar performance as a cross-attention variant. We also introduce a pixel-aligned confidence measure to further guide the diffusion model in uncertain regions with missing details and artifacts. Our experiments show that GScenes outperforms existing pose-free reconstruction techniques in scene reconstruction and performs competitively with state-of-the-art posed sparse-view reconstruction methods.
@@ Appendix B Related Work
Both NeRF and 3DGS rely on hundreds of scene captures for photorealistic novel view synthesis. When the input set becomes sparse, the problem becomes ill-posed, as several simultaneous 3D representations can agree with the training set. Regularization techniques are amongst the earliest techniques to address this limitation. Typical methods leverage depth from Structure-from-Motion (SfM)(Deng et al., 2022; Roessle et al., 2022), monocular estimation(Li et al., 2024; Xiong et al., 2023; Zhu et al., 2023; Chung et al., 2023), or RGB-D sensors (Wang et al., 2023a). DietNeRF (Jain et al., 2021) uses a semantic consistency loss based on CLIP (Radford et al., 2021) features, while FreeNeRF (Yang et al., 2023) regularizes the frequency range of NeRF inputs. RegNeRF (Niemeyer et al., 2022) and DiffusioNeRF (Wynn & Turmukhambetov, 2023) maximize the likelihoods of rendered patches using normalizing flows or diffusion models, respectively. However, such techniques usually fail under extreme sparsity like 3, 6, or 9 input images for a $360^{\circ}$ scene due to weaker priors. Generative priors can be viewed as a stronger form of regularization as they provide extrapolation capabilities for inferring details in unknown parts of a scene.
@@
When only a few or a single view is available, regularization techniques are often insufficient to resolve reconstruction ambiguities. To address this, recent research focuses on training priors for novel view synthesis across multiple scenes. pixelNeRF (Yu et al., 2021) uses pixel-aligned CNN features as conditioning for a shared NeRF MLP, while other approaches (Trevithick & Yang, 2021; Chen et al., 2021; Henzler et al., 2021; Lin et al., 2023b) condition NeRF on 2D or fused 3D features. Further priors have been learned on triplanes (Irshad et al., 2023), voxel grids (Guo et al., 2022), and neural points (Wewer et al., 2023). Leveraging 3D Gaussian Splatting (Kerbl et al., 2023), methods like pixelSplat (Charatan et al., 2024) and MVSplat (Chen et al., 2024) achieve state-of-the-art performance in stereo view interpolation. However, these regression-based techniques infer blurry novel views in case of high uncertainty.
@@
While diffusion models have been applied directly on 3D representations like triplanes (Shue et al., 2023; Chen et al., 2023a), voxel grids (Müller et al., 2023), or (neural) point clouds (Zhou et al., 2021; Melas-Kyriazi et al., 2023; Schröppel et al., 2024), 3D data is scarce. Given the success of large-scale diffusion models for image synthesis, there is a great research interest in leveraging them as priors for 3D reconstruction and generation. DreamFusion (Poole et al., 2023) and follow-ups (Wang et al., 2023b; Lin et al., 2023a; Chen et al., 2023b; Deng et al., 2023; Tang et al., 2024) employ score distillation sampling (SDS) to iteratively maximize the likelihood of radiance field renderings under a conditional 2D diffusion prior. For sparse-view reconstruction, existing approaches incorporate view-conditioning via epipolar feature transform (Zhou & Tulsiani, 2023), cross-attention to encoded relative poses (Liu et al., 2023; Sargent et al., 2024), or pixelNeRF (Yu et al., 2021) feature renderings (Wu et al., 2024). However, this fine-tuning is expensive and requires large-scale multi-view data, which we circumvent with GScenes.
@@
For building a generalizable sparse-view reconstruction method, the assumption of camera poses during inference limits applications to real-world scenarios where usually only an uncalibrated set of 2D images are available with no known camera extrinsics or intrinsics. Several recent works (Jiang et al., 2024; Fan et al., 2024) have attempted to solve reconstruction in a pose-free setting by jointly optimizing poses and NeRF or 3D Gaussian parameters during scene optimization. These methods typically outperform previous techniques (Chen & Lee, 2023; Lin et al., 2021; Bian et al., 2023; Fu et al., 2023) where reconstruction is done in two stages - first estimating poses and then optimizing the 3D representation. However, errors in the initial pose estimation harm subsequent scene optimization, resulting in inferior NVS quality. In our work, we use the MASt3R pipeline with 3DGS for predicting 3D Gaussians and camera parameters in a global coordinate system from a set of unposed 2D images.

### v3

@@ Abstract
In this work, we introduce a generative approach for pose-free (without camera parameters) reconstruction of $360^{\circ}$scenes from a sparse set of 2D images. Pose-free scene reconstruction from incomplete, pose-free observations is usually regularized with depth estimation or 3D foundational priors. While recent advances have enabled sparse-view reconstruction of large complex scenes (with high degree of foreground and background detail) with known camera poses using view-conditioned generative priors, these methods cannot be directly adapted for the pose-free setting when ground-truth poses are not available during evaluation. To address this, we propose an image-to-image generative model designed to inpaint missing details and remove artifacts in novel view renders and depth maps of a 3D scene. We introduce context and geometry conditioning using Feature-wise Linear Modulation (FiLM) modulation layers as a lightweight alternative to cross-attention and also propose a novel confidence measure for 3D Gaussian splat representations to allow for better detection of these artifacts. By progressively integrating these novel views in a Gaussian-SLAM-inspired process, we achieve a multi-view-consistent 3D representation. Evaluations on the MipNeRF360 and DL3DV-10K benchmark dataset demonstrate that our method surpasses existing pose-free techniques and performs competitively with state-of-the-art posed (precomputed camera parameters are given) reconstruction methods in complex $360^{\circ}$scenes. Our project page [^1] https://gaussianscenes.github.io provides additional results, videos, and code.
@@ 1 Introduction
Although most previous sparse view 3D scene reconstruction works are posed and require ground-truth camera poses (DSNeRF (Deng et al., 2022), DNGaussian (Li et al., 2024), SparseGS (Xiong et al., 2023), SparseNeRF (Wang et al., 2023a), DietNeRF (Jain et al., 2021), FreeNeRF (Yang et al., 2023), RegNeRF (Niemeyer et al., 2022), DiffusioNeRF (Wynn and Turmukhambetov, 2023)), recent works like COGS (Jiang et al., 2024) and InstantSplat (Fan et al., 2024) have found initial success in the pose-free sparse view reconstruction problem. Although they have been successful in reducing 3D reconstruction-related artifacts such as blur, floaters, color, and streak-like artifacts, they lack generative capabilities for a complete $360^{\circ}$reconstruction, which requires robust priors from powerful generative models (Saharia et al., 2022; Rombach et al., 2022). In addition, most of these methods lack mechanisms to robustly identify and locate these artifacts in the reconstructed scenes. Recent methods like ZeroNVS (Sargent et al., 2024), Reconfusion (Wu et al., 2024), and CAT3D (Gao* et al., 2024) incorporate 3D view conditioning for realistic extrapolation but depend on accurate poses and cannot be trivially extended to the pose-free problem. Gaussian Object (Yang et al., 2024), iFusion (Wu et al., 2023), and UpFusion (Nagoor Kani et al., 2024) do provide pose-free generative solutions but remain limited to only object reconstruction rather than scene reconstruction. Other generative solutions like ZeroNVS, ReconFusion, and CAT3D use stronger priors over regularization-based techniques like FreeNeRF, RegNeRF, and DietNeRF but rely on large-scale 3D datasets, video datasets, and compute resources not available to the average researcher. Furthermore, works like ReconFusion and CAT3D remain closed-source with no access to their models, data or other reproducibility data, which limits their usage for the community. We summarize such existing methods and their use cases and attributes in Fig 2.
@@ 2 Method
This section begins with an overview of our method in Sec. 2.1, detailing our approach for reconstructing a 3D scene from a sparse set of uncalibrated 2D images. In Sec. 2.1, we describe how we initialize a Gaussian point cloud using MASt3R and 3DGS to provide a coarse 3D representation. Sec. 2.2 introduces our RGBD image-to-image generative model, which refines rendered novel views by correcting artifacts and filling in missing details. We propose a confidence measure (Sec. 2.3) based on cumulative transmittance and Gaussian density to guide the generative model toward unreliable regions. Sec. 2.4 outlines our synthetic dataset creation process for training with high-quality RGBD supervision, followed by depth-augmented autoencoder finetuning (Sec. 2.5) and UNet finetuning with synthesized data (Sec. 2.5.1) to improve generation quality. Finally, Secs. 2.5.2 and 2.6 detail the inference pipeline and 3D Gaussian optimization process, and Sec. 2.7 describes a test-time pose refinement step.
@@ 2.1 Algorithm Overview
Given a set of $M$ images $\mathcal{I}=\{I_{1},I_{2},\dots,I_{M}\}$ of an underlying 3D scene with unknown intrinsics and extrinsics, our goal is to reconstruct the 3D scene, estimate the camera poses of a monocular camera at the $M$ training views, and synthesize novel views at evaluation time given by $N$ unseen test images $\{I_{M+1},I_{M+2},\dots,I_{M+N}\}$.
@@
1:Sparse image set $\mathcal{I}=\{I_{1},\dots,I_{M}\}$, densification interval $k$, iterations NN
@@
[…]
@@
8:   $(\hat{I},\hat{D})\leftarrow R_{\pi_{novel}}(\mathcal{G})$ $\triangleright$ Render image and depth map
@@
10:   Extract semantic context & geometry embeddings $\{c_{clip},c_{geo}\}$ $\triangleright$ Sec 2.2.2
11:   $(I_{ref},D_{ref})\leftarrow\text{Refine}(\hat{I},\hat{D},\mathcal{C},c_{clip},c_{geo})$ $\triangleright$ Refine $\hat{I}$, $\hat{D}$ with generative priors (Sec 2.5.2)
[…]
@@
16:  Optimize 3D Gaussians $\mathcal{G}$ $\triangleright$ Eq. 6 or 3DGS loss
@@ 2.2.1 Generative Model Architecture
Our model comprises (1) A variational autoencoder (encoder $\mathcal{E}$, decoder $\mathcal{D}$) operating in a compressed latent space. (2) A UNet denoiser $\epsilon_{\theta}$ predicting noise in diffused latent $z_{t}$ (3) A multi-modal conditioning incorporating RGBD renders, confidence maps, semantic context, and geometric information.
@@ 2.2.2 Multi-modal Conditioning
[…]
@@ 2.2.3 Parameter-Efficient FiLM Conditioning
[…]
@@ 2.3 Pixel-Aligned Confidence Map
Figure 7: Confidence Measure comparison with 3DGS-Enhancer (Liu et al., 2024). Our confidence map accurately identifies artifacts and 0-Gaussian regions in the sparse-view (darker pixels) while  Liu et al. (2024) incorrectly attributes high confidence to regions with overlap of small-scale Gaussians. NVS render from a densely fitted 3DGS representation is provided for reference.
@@ 2.4 RGBD Dataset Creation
For training the additional weights in $\bm{\epsilon_{\theta}}$ for RGBD image-to-image diffusion, we rely on a set $\mathcal{X}=\{(I^{i},\hat{I}^{i},\mathcal{C}^{i},c_{clip}^{i},c_{geo}^{i})_{i=1}^{N}\}$, each containing a clean RGBD image $I^{i}$, an RGBD image with artifacts $\hat{I}^{i}$ and the corresponding confidence map $\mathcal{C}^{i}$, CLIP features of source images $c_{clip}^{i}\in\mathcal{R}^{M\times 768}$, and plücker embeddings of source and target cameras $c_{geo}^{i}\in\mathcal{R}^{(M+1)\times 78}$, to “teach” the diffusion model how to inpaint missing details and detect Gaussian artifacts guided by the confidence map, context and geometry features and generate a clean version of the conditioning image. For this, we build a dataset generation pipeline comprising a high-quality 3DGS model fitted to dense views, a low-quality 3DGS model fitted to few views, and camera interpolation and perturbation modules to use supervision of the high-quality model at viewpoints beyond ground truth camera poses. For a given scene, we fit sparse models for $M\in\{3,6,9,18\}$ number of views. We render $I^{i}$ using the high-quality model and $\hat{I}^{i}$, $\mathcal{C}^{i}$ using the low-quality model. We save the CLIP features $c_{clip}$ and plücker embeddings of the $M$ sparse views and 1 target view per sample for conditioning $\bm{\epsilon_{\theta}}$ . Our fine-tuning setup is illustrated in Fig  12 (Sec C).
@@ 2.6 Scene Reconstruction with Generative Priors
Our generative diffusion model provides a generative prior to infer plausible detail in unobserved regions. Despite view conditioning using pose embeddings, the generated images at novel poses lack complete 3D consistency. For this, we devise an iterative strategy where we first sample novel views along an elliptical trajectory fitted to the training views. We initialize the Gaussian optimization with the set of Gaussians $\mathcal{G}$ fitted to the training views (Sec 2.1). Each novel view is added to the training stack at the beginning of every densification step to encourage the optimization to adjust to the distilled scene priors. At every iteration, we sample either an observed or unobserved viewpoint from the current training stack. We bring back Adaptive Density Control to encourage densification of Gaussians in 0-Gaussian regions. We employ the 3DGS objective for the training views. For novel views, we employ the SparseFusion (Zhou and Tulsiani, 2023) objective in the RGB space and a PCC loss for the rendered and denoised depths.
@@ 2.7 Test-time Pose Alignment
MASt3R + 3DGS RegNeRF ´ DiffusioNeRF ´ ZeroNVS* ´ COGS Ours Ground Truth
@@
MASt3R + 3DGS ZeroNVS* ´ DiffusioNeRF ´ COGS Ours Ground Truth
@@ 3.3 Comparative Results
MipNeRF360 FreeNeRF ´ 11.886 12.874 13.673 0.180 0.218 0.236 0.675 0.654 0.638
RegNeRF ´ 12.294 13.204 13.796 0.182 0.207 0.217 0.668 0.656 0.625
DiffusioNeRF ´ 9.996 12.025 12.567 0.159 0.211 0.213 0.715 0.648 0.659
ZeroNVS* ´ 11.991 11.817 11.729 0.142 0.129 0.128 0.710 0.705 0.694
ReconFusion ´ 15.50 16.93 18.19 0.358 0.401 0.432 0.585 0.544 0.511
CAT3D ´ 16.62 17.72 18.67 0.377 0.425 0.460 0.515 0.482 0.460
MASt3R + 3DGS 12.585 14.285 15.042 0.231 0.279 0.310 0.593 0.550 0.531
@@
MipNeRF360 FreeNeRF ´ 354.244 353.480 346.874 0.261 0.271 0.282 0.388 0.369 0.365
RegNeRF ´ 358.855 360.380 338.183 0.281 0.301 0.264 0.407 0.405 0.377
DiffusioNeRF ´ 370.346 347.522 342.373 0.293 0.267 0.257 0.441 0.386 0.402
ZeroNVS* ´ 356.395 350.920 343.930 0.283 0.296 0.300 0.433 0.426 0.413
MASt3R + 3DGS 294.819 252.830 231.169 0.210 0.164 0.133 0.303 0.273 0.264
@@ 4 Conclusion
In this work, we present GScenes where we integrate an image-to-image RGBD diffusion model with a pose-free reconstruction pipeline in MASt3R to reconstruct a $360^{\circ}$3D scene from a few uncalibrated 2D images. We introduce context and geometry conditioning through FiLM layers, achieving similar performance as a cross-attention variant. We also introduce a pixel-aligned confidence measure to further guide the diffusion model in uncertain regions with missing details and artifacts. Our experiments show that GScenes outperforms existing pose-free reconstruction techniques in scene reconstruction and performs competitively with state-of-the-art posed sparse-view reconstruction methods.
5 Acknowledgement
This work was supported by Army Research Laboratory award W911NF2320008 and ONR with N00014-21-1-2812.
@@ Appendix B Related Work
Both NeRF and 3DGS rely on hundreds of scene captures for photorealistic novel view synthesis. When the input set becomes sparse, the problem becomes ill-posed, as several simultaneous 3D representations can agree with the training set. Regularization techniques are amongst the earliest techniques to address this limitation. Typical methods leverage depth from Structure-from-Motion (SfM)(Deng et al., 2022; Roessle et al., 2022), monocular estimation(Li et al., 2024; Xiong et al., 2023; Zhu et al., 2023; Chung et al., 2023), or RGB-D sensors (Wang et al., 2023a). DietNeRF (Jain et al., 2021) uses a semantic consistency loss based on CLIP (Radford et al., 2021) features, while FreeNeRF (Yang et al., 2023) regularizes the frequency range of NeRF inputs. RegNeRF (Niemeyer et al., 2022) and DiffusioNeRF (Wynn and Turmukhambetov, 2023) maximize the likelihoods of rendered patches using normalizing flows or diffusion models, respectively. However, such techniques usually fail under extreme sparsity like 3, 6, or 9 input images for a $360^{\circ}$ scene due to weaker priors. Generative priors can be viewed as a stronger form of regularization as they provide extrapolation capabilities for inferring details in unknown parts of a scene.
@@
When only a few or a single view is available, regularization techniques are often insufficient to resolve reconstruction ambiguities. To address this, recent research focuses on training priors for novel view synthesis across multiple scenes. pixelNeRF (Yu et al., 2021) uses pixel-aligned CNN features as conditioning for a shared NeRF MLP, while other approaches (Trevithick and Yang, 2021; Chen et al., 2021; Henzler et al., 2021; Lin et al., 2023b; Szymanowicz et al., 2023) condition NeRF on 2D or fused 3D features. Further priors have been learned on triplanes (Irshad et al., 2023), voxel grids (Guo et al., 2022), neural points (Wewer et al., 2023), and IB-planes Anciukevičius et al. (2024). Leveraging 3D Gaussian Splatting (Kerbl et al., 2023), methods like pixelSplat (Charatan et al., 2024) and MVSplat (Chen et al., 2024) achieve state-of-the-art performance in stereo view interpolation. However, these regression-based techniques infer blurry novel views in case of high uncertainty.
@@
While diffusion models have been applied directly on 3D representations like triplanes (Shue et al., 2023; Chen et al., 2023a), voxel grids (Müller et al., 2023), or (neural) point clouds (Zhou et al., 2021; Melas-Kyriazi et al., 2023; Schröppel et al., 2024), 3D data is scarce. Given the success of large-scale diffusion models for image synthesis, there is a great research interest in leveraging them as priors for 3D reconstruction and generation. DreamFusion (Poole et al., 2023) and follow-ups (Wang et al., 2023b; Lin et al., 2023a; Chen et al., 2023b; Deng et al., 2023; Tang et al., 2024) employ score distillation sampling (SDS) to iteratively maximize the likelihood of radiance field renderings under a conditional 2D diffusion prior. For sparse-view reconstruction, existing approaches incorporate view-conditioning via epipolar feature transform (Zhou and Tulsiani, 2023), cross-attention to encoded relative poses (Liu et al., 2023; Sargent et al., 2024), or pixelNeRF (Yu et al., 2021) feature renderings (Wu et al., 2024). However, this fine-tuning is expensive and requires large-scale multi-view data, which we circumvent with GScenes.
@@
For building a generalizable sparse-view reconstruction method, the assumption of camera poses during inference limits applications to real-world scenarios where usually only an uncalibrated set of 2D images are available with no known camera extrinsics or intrinsics. Several recent works (Jiang et al., 2024; Fan et al., 2024) have attempted to solve reconstruction in a pose-free setting by jointly optimizing poses and NeRF or 3D Gaussian parameters during scene optimization. These methods typically outperform previous techniques (Chen and Lee, 2023; Lin et al., 2021; Bian et al., 2023; Fu et al., 2023) where reconstruction is done in two stages - first estimating poses and then optimizing the 3D representation. However, errors in the initial pose estimation harm subsequent scene optimization, resulting in inferior NVS quality. In our work, we use the MASt3R pipeline with 3DGS for predicting 3D Gaussians and camera parameters in a global coordinate system from a set of unposed 2D images.